OpenStudy (anonymous):

Parametric Differentiation x= 2+ sec theta , y = 1+ 2tan theta theta = pi/6 (30 degrees) Find dy/dx and evaluate the specified value of the parameter

OpenStudy (anonymous):

so first i differentiated and got \[2\sec ^{2}\theta/\sec \theta \tan \theta\]

OpenStudy (amistre64):

<sec(t)tan(t),2sec^2(t)>

OpenStudy (anonymous):

then i substituted the value in and got 3

OpenStudy (anonymous):

is 3 correct?

OpenStudy (amistre64):

maybe, gotta calc it :)

OpenStudy (anonymous):

\[((2)(\sqrt{3}/2))/(\sqrt{3}/2*1/\sqrt{3})\]

OpenStudy (amistre64):

i get 1/4

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

i think i know where i went wrong let me check it

OpenStudy (amistre64):

\[(\frac{2}{\sqrt{3}})^2 \div \frac{2}{\sqrt{3}}\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

the top maybe wrong i got the same bottom when i did it over

OpenStudy (amistre64):

dropped a '2' :) \[\frac{2\over3}{8\over 3}={2\over8}=1/4\]

OpenStudy (anonymous):

give me a sec to figure it out

OpenStudy (anonymous):

i got 8/2

OpenStudy (anonymous):

for the top part it was 2 sec^2 theta

OpenStudy (amistre64):

youre right, i got it flipped in me head

OpenStudy (anonymous):

so its 4? yay!

OpenStudy (amistre64):

its 4 :)

OpenStudy (anonymous):

as always thank youso much for your help amistre

OpenStudy (amistre64):

thanx for correcting my mistake ;)

OpenStudy (anonymous):

your welcomes!